pearis:

Very Important please help!! Using the following equation, find the center and radius of the circle. You must show all work and calculations to receive credit. x2 −2x + y2 − 6y = 26

ZombieLover:

Given: x^2 −2x + y^2 − 6y = 26 complete the square: x^2 -2x + (-2/2)^2 + y^2 - 6y + (-6/2)^2 = 26 + (-2/2)^2 + (-6/2)^2 (x - 1)^2 + (y - 3)^2 = 26 + 1 + 9 (x - 1)^2 + (y - 3)^2 = 36 (x - 1)^2 + (y - 3)^2 = 6^2 center: (1,3) radius: 6

pearis:

thank you so much!!

ZombieLover:

Can you close the question?